OpenStudy (anonymous):

Write -2x^2(-5x^2+4x^3) in standard form

OpenStudy (anonymous):

Hm??

OpenStudy (marvin456):

Go with the first link that one seems better.

OpenStudy (anonymous):

-156

OpenStudy (misty1212):

HI!

OpenStudy (misty1212):

\[ -2x^2(-5x^2+4x^3) \] distribute first step is \[-2x^2\times (-5x^2)+(-2x^2)\times (4x^3)\]

OpenStudy (marvin456):

Hi misty! :)

OpenStudy (misty1212):

lol the answer is not a number!

OpenStudy (anonymous):

-2x4 * (4x - 5)

OpenStudy (anonymous):

Step 1 : Multiply 2x2 by 4x3-5x2

OpenStudy (anonymous):

step 2 Pull out like factors : 4x3 - 5x2 = x2 * (4x - 5)

OpenStudy (misty1212):

oh my this is not so hard

OpenStudy (anonymous):

-8x + 10x

OpenStudy (misty1212):

\[-2x^2\times (-5x^2)+(-2x^2)\times (4x^3)\\ =10x^4-8x^5\]

OpenStudy (misty1212):

standard form will be \[-8x^5-10x^4\]

OpenStudy (misty1212):

@-EtherealEbullience where are your exponents?

OpenStudy (anonymous):

Ah, I see, thank you.

OpenStudy (anonymous):

step 3: Multiplying exponential expressions : x^2 multiplied by x^2 = x^(2 + 2) = x^4

OpenStudy (anonymous):

I didn't know I was suppose to multiply the exponents. x3

OpenStudy (anonymous):

yea then u should have got this 0 - 2x^4 * (4x - 5)

OpenStudy (anonymous):

Now Simplify 0 - 2x^4^(4x-5)

OpenStudy (anonymous):

this is the final step

OpenStudy (anonymous):

-2x^4 * (4x - 5)

OpenStudy (anonymous):

Can i get a medal